OpenStudy (anonymous):

Find the value of this number. 5!

OpenStudy (anonymous):

p e n i s!

OpenStudy (whpalmer4):

n! is (n)(n-1)(n-2)...(2)(1)

OpenStudy (whpalmer4):

Another way of thinking of n! is n * (n-1)!, with 1! = 1. So 2! = 2(1!) = 2*1. 3! = 3*2! = 3*2*1 = 6. 4! = 4*3! = 4*(3*2!) = 4*(3*(2*1!)) = 4*3*2*1, etc.